# A microstructural model of transversely isotropic, fibre-reinforced hydrogels

### v1

@@ Abstract
Fibre-reinforced hydrogels are promising materials for biomedical applications due to their strength, toughness, and tunability. However, it remains unclear how to design fibre-reinforced hydrogels for use in specific applications due to the lack of a flexible modelling framework that can predict and hence optimise their behaviour. In this paper, we present a microstructural model for transversely isotropic fibre-reinforced hydrogels that captures the specific geometry of the fibre network. The model also accounts for slack in the initial fibre network that is gradually removed upon deformation. The mechanical model for the fibre network is coupled to a nonlinear poroelastic model for the hydrogel matrix that accounts for osmotic stress. By comparing the model predictions to data from load-controlled, unconfined compression experiments, we show that the model can capture J-shaped stress-strain curves and time-dependent creep responses. We showcase how the model can be used to guide the design of materials for artificial cartilage by exploring how to maximise interstitial fluid pressure. We find that fluid pressurisation can be increased by using stiffer fibres, removing slack from the fibre network, and reducing the Young’s modulus of the hydrogel matrix. Finally, an open-source Python package has been developed for simulating unconfined compression experiments using the model.
@@ 1 Introduction
Despite the advantages of fibre-reinforced hydrogels, there remain challenges in engineering these materials for use in specific applications. In particular, there is a lack of understanding of how the fibre properties and the geometry of the fibre network should be varied in order to produce tailored macroscropic responses [19]. Mathematical and computational models of varying complexity have been used to shed light on the macroscopic response of fibre-reinforced hydrogels. The Halpin–Tsai (HT) equations [20] are a relatively simple micromechanical model that describes the stiffness of a fibre-reinforced matrix in terms of the fibre and matrix properties. Jordan et al. [15] found that the HT equations do not always provide accurate predictions of the stiffness of fibre-reinforced hydrogels, and attributed discrepancies to the large contrast in the Young’s moduli of the fibres and hydrogel matrix. Beckett et al. [19] also showed that the HT equations have limited applicability to fibre-reinforced hydrogels and proposed that swelling of the hydrogel matrix could invalidate the assumptions that underpin the HT equations. Visser et al. [14] developed a simple expression for the stiffness of transversely isotropic fibre-reinforced hydrogels in terms of the fibre properties. Subsequent studies by Castilho et al. [18] and Chen et al. [21] used three-dimensional (3D) finite element simulations to understand, respectively, the nonlinear elastic and linear poroelastic response of the composites produced by Visser et al. Due to the specific geometry of the fibre network used by Visser et al, in which the fibres are laid down to form a two-dimensional (2D) grid, the modelling results of Castilho, Chen, et al. might not carry over to other systems. Motivated by problems in cartilage mechanics, Ateshian et al. [22, 23] developed a general approach for modelling fibre-reinforced soft hydrated materials. However, simulating these models required the development of a bespoke 3D nonlinear finite element code. The high computational cost of 3D finite element simulations, along with the effort required to create bespoke code, limits their use as predictive design tools.
@@
The purpose of this paper is, therefore, to present a flexible and computationally efficient microstructural modelling framework for fibre-reinforced hydrogels that is implemented in an open-source Python package. The framework links the fibre properties and fibre network geometry to the macroscopic material response. We focus on modelling a class of transversely isotropic materials developed by Moore et al. [16] to mimic the structure and mechanics of articular cartilage. These materials consisted of horizontally aligned fibre networks layered vertically and embedded in a hydrogel matrix, as shown in Fig. 1. Similar composites have been fabricated by a number of other authors [24, 14, 17, 19]. We develop a microstructural model of the hydrogel and fibre phases, explicitly modelling the hydration of the gel, before considering unconfined compression tests of the composite material. We build upon techniques used for modelling fibre networks, and incorporate them within a nonlinear poroelastic framework that accounts for osmotic effects. The poroelastic modelling of the hydrogel component of the material follows standard approaches [25, 26]; hence, the main focus of this paper is on modelling the fibre network.
@@
Figure 1: (a) Photograph (scale bar = 5 mm) of a transversely isotropic fibre-reinforced hydrogel fabricated by Moore et al. [16] along with a magnified 3D reconstruction (scale bar = 50 $\mu$m). Image adapted from Ref. [16]. (b) A mathematical idealisation of a transversely isotropic fibre-reinforced hydrogel, with the layers of fibre network shown as parallel black lines in the cross section.
@@ 2.1 Microstructural modelling of the fibrous phase of the material
We first derive two microstructural constitutive models for the fibrous phase of the material. We consider a two-dimensional network of fibres and assume that it deforms according to an affine mapping (an assumption that is appropriate for high-density fibre networks [44]). Based upon this assumption, we calculate the strain energy in an arbitrary fibre segment with a specified initial angle, and then calculate the total strain energy in the network by integrating this expression over all angles. For the first constitutive model, we assume the fibres are all initially straight, in the second, we assume that they are wavy and are gradually recruited with increasing tensile strain.
@@ 2.1.2 The strain energy function for fibre networks with initially wavy fibres
We find that it is sometimes more computationally efficient to use a semi-analytical approach where the inner integral in (19) is computed exactly and the outer integral is numerically approximated using the trapezoidal rule, which is exponentially accurate due to the periodicity of the integrand [45]. This hybrid approach allows a variety of fibre recruitment stretch distributions $f(\lambda_{c})$ to be used in the model (e.g. uniform, normal, beta, gamma, exponential) when an exact expression for the double integral in (19) is not available. However, in the implementation discussed below, we stick to using the piecewise quartic function defined in equation (20).
@@ 2.3 Material hydration
where $\Omega_{w}=18\times 10^{-6}$ m3 mol-1 is the molar volume of water, $R_{g}=8.314$ J mol-1 K-1 is the universal gas constant, $T=296$ K is temperature, and $\chi$ is the Flory interaction parameter, which characterises the strength of unfavourable interactions between water molecules and solid components (polymers and fibres). If $\chi\gg 1$, then it is energetically costly to hydrate the material; in this case, the amount of water uptake during the hydration phase will be small. The dependence of $\chi$ on the nominal fibre fraction $\Phi_{f}$ reflects the differing affinity of water with the fibre and matrix phases. Its functional form will be discussed in Section 4.1.
@@ 2.4 Balance laws and bulk equations
Having defined the free energy density of the material, $W=W_{e}+W_{\text{mix}}$, the constitutive relations for a three-dimensional, time-dependent model can now be derived using non-equilibrium thermodynamics. For brevity, we only present the most important equations here; their derivations are provided in Appendix B. We choose to formulate the governing equations using a reference configuration associated with a freely swollen, hydrated state (to be determined in Section 4). Here, freely swollen means that there are no geometric constraints or loads applied to the material. Thus, we let the deformation gradient, $\boldsymbol{\mathsf{F}}_{h}$, describe how the non-hydrated material deforms upon hydration; see Fig. 3. The volume change due to hydration is given by $J_{h}=\det\boldsymbol{\mathsf{F}}_{h}$. We assume that the volume of the hydrated material, $V_{h}$, is equal to the sum of the initial volume of the solid phase, $V_{s}$, and the added volume of water, $V_{w}$. Therefore, $V_{h}=V_{s}+V_{w}$ and $J_{h}=V_{h}/V_{s}=1/(1-\phi_{0})$, where $\phi_{0}=V_{w}/V_{h}$ is the initial volume fraction of water in the hydrated state. That is, $\phi_{0}$ represents the initial porosity of the material.
Figure 3: The three spatial configurations used in the model and the deformation gradient tensors used to move between them. The initial (reference) hydrated state is used as a reference configuration to formulate the bulk equations. The current (deformed) state corresponds to an unconfined compression experiment, where a cylindrical sample has been placed between two plattens and an axial force, $F$, has been applied.
Subsequent deformations relative to the hydrated state are described by the deformation gradient tensor $\boldsymbol{\mathsf{F}}$. By introducing the displacement relative to the hydrated state, $\boldsymbol{u}$, the deformation gradient tensor can be written as $\boldsymbol{\mathsf{F}}=\boldsymbol{\mathsf{I}}+\nabla\boldsymbol{u}$, where $\boldsymbol{\mathsf{I}}$ is the identity tensor and $\nabla$ is the gradient with respect to coordinates in the hydrated state. Similarly, $J=\det\boldsymbol{\mathsf{F}}$ describes how material elements change their volume relative to the hydrated state. The total deformation gradient tensor that describes the deformation from the non-hydrated configuration to the current configuration is given by $\boldsymbol{\mathsf{F}}_{t}=\boldsymbol{\mathsf{F}}\boldsymbol{\mathsf{F}}_{h}$. The total volume change relative to the non-hydrated state is $J_{t}=\det\boldsymbol{\mathsf{F}}_{t}=JJ_{h}$.
@@ 4 Hydration of a fibre-reinforced hydrogel
where we have used $\boldsymbol{\mathsf{F}}=\boldsymbol{\mathsf{I}}$ for the hydrated state and we recall from Sec. 2.4 that $\phi_{0}=1-1/J_{h}$ is the volume fraction of fluid in the initial hydrated state. Thus, in a free-swelling scenario, the elastic and osmotic stresses must balance. Moreover, it follows that the osmotic stress is tensile and acts to stretch the material.
@@
where $\alpha_{\parallel}$ and $\alpha_{\perp}$ are stretches in the direction parallel and perpendicular to the material’s plane of isotropy. This leads to an elastic stress tensor of the form
@@ 4.1 Case study: swelling of PCL-reinforced gelatin
We first fix the nominal volume fraction of fibres to $\Phi_{f}=0.26$ and vary the Flory interaction parameter $\chi$. We consider fibre networks that are initially fully engaged with $\lambda_{m}=1$ and partially engaged with $\lambda_{m}=3$ (recall that $\lambda_{m}$ is the maximum recruitment stretch in the distibution function (20)). In both cases, the deformation due to hydration is strongly anisotropic, with the radial stretch being much smaller than the axial stretch; see Fig. 4 (a)–(b). The anisotropic swelling is caused by the large contrast between the Young’s moduli of the fibres and the matrix. As the Flory interaction parameter increases, the radial and axial stretches decrease because the volume of water absorbed by the matrix decreases, which is reflected in the porosity of the hydrated state strongly decreasing with $\chi$ (Fig. 4 (c)). As the Flory interaction parameter increases, it becomes less energetically favourable to swell the network and hence the osmotic stress decreases (Fig. 4 (d)). In the case of a fully engaged fibre network ($\lambda_{m}=1$), there is virtually no in-plane expansion of the material and all of the deformation is perpendicular to the plane of the fibres. The small in-plane deformation is expected because the osmotic stress responsible for hydration, $\Pi\sim 10$ kPa, is much smaller than the in-plane Young’s modulus, which is dominated by the fibre modulus $E_{f}\sim 400$ MPa. Conversely, the out-of-plane Young’s modulus is controlled by the gelatin matrix, $E_{m}\sim 4$ kPa, and is comparable to the osmotic stress; hence, there is a substantial out-of-plane deformation. When the fibre network is only partially engaged ($\lambda_{m}=3$), we see there is a larger in-plane deformation, which is due to initially inactive fibres causing an effective reduction in the in-plane stiffness of the material, and a smaller out-of-plane deformation. However, the in-plane stretching caused by hydration is always less than the maximum recruitment stretch. Therefore, hydration is not sufficient to fully engage the fibre network. Remarkably, the out-of-plane deformation decreases in such a way that the porosity of the material remains approximately the same in both cases. Thus, the degree of hydration is not significantly impacted by the initial level of fibre engagement.
Figure 4: Hydration of fibre-reinforced hydrogels without (solid lines) and with (dashed lines) recruitment. The panels show the in-plane stretch (a, e), out-of-plane stretch (b, f), volume change due to hydration (c, g), and fluid pressure (d, h) as a function of the Flory interaction parameter, $\chi$, with $\Phi_{f}=0.26$ (a-d), and the nominal fibre fraction, $\Phi_{f}$, with $\chi=0.57$ (e-h).
@@
By fixing $\chi=0.57$ and considering fibre networks that are initially fully engaged ($\lambda_{m}=1$), we see that the deformation and porosity do not strongly depend on the nominal fibre fraction (Fig. 4 (e)–(g)). The osmotic stress decreases linearly with $\Phi_{f}$ (Fig. 4 (h)), reflecting the balance between osmotic stress and the out-of-plane elastic stress produced by the gelatin matrix, the latter of which is proportional to $1-\Phi_{f}$; see Appendix C. When initially inactive fibres are accounted for ($\lambda_{m}=3$), the hydrated state leads to greater in-plane stretching and reduced out-of-plane stretching compared to the active case (as is the case when these parameters are plotted as a function of $\chi$). Moreover, both the in-plane and out-of-plane stretches now strongly depend on the fibre fraction $\Phi_{f}$. The dependence of the stretches on $\Phi_{f}$ ultimately results from the in-plane softening that occurs due to inactive fibres. This softening allows the osmotic stress to drive a greater in-plane expansion of the material. As the nominal fibre fraction increases, the in-plane stiffness increases as well, reducing the in-plane deformation and increasing the out-of-plane deformation. Again, whether the fibres are initially active or inactive does not affect the porosity as a function of $\Phi_{f}$ (Fig. 4 (g)).

### v2

@@ Abstract
Fibre-reinforced hydrogels are promising materials for biomedical applications due to their strength, toughness, and tunability. However, it remains unclear how to design fibre-reinforced hydrogels for use in specific applications due to the lack of a flexible modelling framework that can predict and hence optimise their behaviour. In this paper, we present a microstructural model for transversely isotropic fibre-reinforced hydrogels that captures the specific geometry of the fibre network. The model also accounts for slack in the initial fibre network that is gradually removed upon deformation. The mechanical model for the fibre network is coupled to a nonlinear poroelastic model for the hydrogel matrix that accounts for osmotic stress. By comparing the model predictions to data from unconfined compression experiments, we show that the model can capture J-shaped stress-strain curves and time-dependent creep responses. We showcase how the model can be used to guide the design of materials for artificial cartilage by exploring how to maximise interstitial fluid pressure. We find that fluid pressurisation can be increased by using stiffer fibres, removing slack from the fibre network, and reducing the Young’s modulus of the hydrogel matrix. Finally, a high-level and open-source Python package has been developed for simulating unconfined compression experiments using the model.
@@ 1 Introduction
Despite the advantages of fibre-reinforced hydrogels, there remain challenges in engineering these materials for use in specific applications. In particular, there is a lack of understanding of how the fibre properties and the geometry of the fibre network should be varied in order to produce tailored macroscropic responses [19]. Mathematical and computational models of varying complexity have been used to shed light on the macroscopic response of fibre-reinforced hydrogels. The Halpin–Tsai (HT) equations [20] are a relatively simple micromechanical model that describes the stiffness of a fibre-reinforced matrix in terms of the fibre and matrix properties. Jordan et al. [15] found that the HT equations do not always provide accurate predictions of the stiffness of fibre-reinforced hydrogels, and they attributed discrepancies to the large contrast in the Young’s moduli of the fibres and the hydrogel matrix. Beckett et al. [19] also showed that the HT equations have limited applicability to fibre-reinforced hydrogels and proposed that swelling of the hydrogel matrix could invalidate the assumptions that underpin the HT equations. Visser et al. [14] developed a simple expression for the stiffness of transversely isotropic fibre-reinforced hydrogels in terms of the fibre properties. Subsequent studies by Castilho et al. [18] and Chen et al. [21] used three-dimensional (3D) finite element simulations to understand, respectively, the nonlinear elastic and linear poroelastic response of the composites produced by Visser et al. Due to the specific geometry of the fibre network used by Visser et al, in which the fibres are laid down to form a two-dimensional (2D) grid, the modelling results of Castilho, Chen, et al. might not carry over to other systems. Motivated by problems in cartilage mechanics, Ateshian et al. [22, 23] developed a general approach for modelling fibre-reinforced soft hydrated materials. However, simulating these models required the development of a bespoke 3D nonlinear finite element code. The high computational cost of 3D finite element simulations, along with the effort required to create bespoke code, reduces their use as predictive design tools.
@@
The purpose of this paper is, therefore, to present a flexible and computationally efficient microstructural modelling framework for fibre-reinforced hydrogels that is implemented in a high-level open-source Python package. The modelling framework links the fibre properties and fibre network geometry to the macroscopic material response. We focus on modelling a class of transversely isotropic materials developed by Moore et al. [16] to mimic the structure and mechanics of articular cartilage. These materials consisted of horizontally aligned fibre networks layered vertically and embedded in a hydrogel matrix, as shown in Fig. 1. Similar composites have been fabricated by a number of other authors [24, 14, 17, 19]. We develop a microstructural model of the hydrogel and fibre phases, explicitly modelling the hydration of the gel, before considering unconfined compression tests of the composite material. We build upon techniques used for modelling fibre networks, and incorporate them within a nonlinear poroelastic framework that accounts for osmotic effects. The poroelastic modelling of the hydrogel component of the material follows standard approaches [25, 26]; hence, the main focus of this paper is on modelling the fibre network.
@@
Figure 1: (a) Photograph (scale bar = 5 mm) of a transversely isotropic fibre-reinforced hydrogel fabricated by Moore et al. [16] along with a 3D reconstruction using nano computed tomography (scale bar = 50 $\mu$m). (b) A mathematical idealisation of a transversely isotropic fibre-reinforced hydrogel, with the layers of fibre network shown as parallel black lines in the cross section.
@@ 2.1 Microstructural modelling of the fibrous phase of the material
We first derive two microstructural constitutive models for the fibrous phase of the material. We consider a two-dimensional network of fibres and assume that it deforms according to an affine mapping (an assumption that is appropriate for high-density fibre networks [44]). Based upon this assumption, we calculate the strain energy in an arbitrary fibre segment with a specified initial angle, and then calculate the total strain energy in the network by integrating this expression over all angles. For the first constitutive model, we assume the fibres are all initially straight; in the second, we assume that they are wavy and are gradually recruited with increasing tensile strain.
@@ 2.1.2 The strain energy function for fibre networks with initially wavy fibres
We find that it is sometimes more computationally efficient to use a semi-analytical approach where the inner integral in (19) is computed exactly and the outer integral is numerically approximated using the trapezoidal rule, which is exponentially accurate due to the periodicity of the integrand [45]. This hybrid approach allows a variety of fibre recruitment stretch distributions $f(\lambda_{c})$ to be used in the model (e.g. uniform, normal, beta, gamma, exponential) when an exact expression for the double integral in (19) is not available. However, in the implementation discussed below, we use the piecewise quartic function defined in equation (20).
@@ 2.3 Material hydration
where $\Omega_{w}=18\times 10^{-6}$ m3 mol-1 is the molar volume of water, $R_{g}=8.314$ J mol-1 K-1 is the universal gas constant, $T=296$ K is temperature, and $\chi$ is the Flory interaction parameter, which characterises the strength of unfavourable interactions between water molecules and solid components (polymers and fibres). If $\chi\gg 1$, then it is energetically costly to hydrate the material; in this case, the amount of water uptake during the hydration phase will be small. The dependence of $\chi$ on the nominal fibre fraction $\Phi_{f}$ reflects the differing affinity of water to the fibre and matrix phases. Its functional form will be discussed in Section 4.1.
@@ 2.4 Balance laws and bulk equations
Having defined the free energy density of the material, $W=W_{e}+W_{\text{mix}}$, the constitutive relations for a three-dimensional, time-dependent model can now be derived using non-equilibrium thermodynamics. For brevity, we only present the most important equations here; their derivations are provided in Appendix B. We choose to formulate the governing equations using a reference configuration associated with a freely swollen, hydrated state (to be determined in Section 4). Here, freely swollen means that there are no geometric constraints or loads applied to the material. Thus, we let the deformation gradient, $\boldsymbol{\mathsf{F}}_{h}$, describe how the non-hydrated material deforms upon hydration; see Fig. 3. The relative volume change due to hydration is given by $J_{h}=\det\boldsymbol{\mathsf{F}}_{h}$. We assume that the volume of the hydrated material, $V_{h}$, is equal to the sum of the initial volume of the solid phase, $V_{s}$, and the added volume of water, $V_{w}$. Therefore, $V_{h}=V_{s}+V_{w}$ and $J_{h}=V_{h}/V_{s}=1/(1-\phi_{0})$, where $\phi_{0}=V_{w}/V_{h}$ is the initial volume fraction of water in the hydrated state. That is, $\phi_{0}$ represents the initial porosity of the material.
Figure 3: The three spatial configurations used in the model and the deformation gradient tensors used to move between them. The unhydrated (nominal) state represents the material as it was prepared before hydration. The hydrated (reference) state corresponds to a freely swollen material and is used as a reference configuration to formulate the model equations. The current (deformed) state in this work corresponds to an unconfined compression experiment, where a cylindrical sample has been placed between two plattens and an axial force, $F$, has been applied.
Subsequent deformations relative to the hydrated state are described by the deformation gradient tensor $\boldsymbol{\mathsf{F}}$. By introducing the displacement relative to the hydrated state, $\boldsymbol{u}$, the deformation gradient tensor can be written as $\boldsymbol{\mathsf{F}}=\boldsymbol{\mathsf{I}}+\nabla\boldsymbol{u}$, where $\boldsymbol{\mathsf{I}}$ is the identity tensor and $\nabla$ is the gradient with respect to coordinates in the hydrated state. Similarly, $J=\det\boldsymbol{\mathsf{F}}$ describes the volume of a material element in the current (deformed) state relative to its volume in the hydrated (reference) state. The total deformation gradient tensor that describes the deformation from the non-hydrated (nominal) configuration to the current (deformed) configuration is given by $\boldsymbol{\mathsf{F}}_{t}=\boldsymbol{\mathsf{F}}\boldsymbol{\mathsf{F}}_{h}$. The total volume change of a material element relative to its volume in the non-hydrated state is $J_{t}=\det\boldsymbol{\mathsf{F}}_{t}=JJ_{h}$.
@@ 4 Hydration of a fibre-reinforced hydrogel
where we have used $\boldsymbol{\mathsf{F}}=\boldsymbol{\mathsf{I}}$ for the hydrated state and we recall from Sec. 2.4 that $\phi_{0}=1-1/J_{h}$ is the volume fraction of fluid in the initial hydrated state. Equation (32) implies that in a free-swelling scenario, the elastic and osmotic stresses must balance. Moreover, since $\Pi(\phi_{0})>0$, it follows that the osmotic stress is tensile and acts to stretch the material.
@@
where $\alpha_{\parallel}$ and $\alpha_{\perp}$ are stretches in the direction parallel and perpendicular to the plane of the fibres, respectively. The form of (36) leads to an elastic stress tensor of the form
@@ 4.1 Case study: swelling of PCL-reinforced gelatin
We first fix the nominal volume fraction of fibres to $\Phi_{f}=0.26$ and vary the Flory interaction parameter $\chi$. We consider fibre networks that are initially fully engaged with $\lambda_{m}=1$ and partially engaged with $\lambda_{m}=3$ (recall that $\lambda_{m}$ is the maximum recruitment stretch in the distribution function (20)). In both cases, the deformation due to hydration is strongly anisotropic, with the radial stretch being much smaller than the axial stretch; see Fig. 4 (a)–(b). The anisotropic swelling is caused by the large contrast between the Young’s moduli of the fibres and the matrix. As the Flory interaction parameter increases, the radial and axial stretches decrease because the volume of water absorbed by the matrix decreases, which is reflected in the porosity of the hydrated state strongly decreasing with $\chi$ (Fig. 4 (c)). Moreover, with increasing $\chi$, it becomes less energetically favourable to swell the network and hence the osmotic stress decreases (Fig. 4 (d)). In the case of a fully engaged fibre network ($\lambda_{m}=1$), there is virtually no in-plane expansion of the material and all of the deformation is perpendicular to the plane of the fibres. The small in-plane deformation is expected because the osmotic stress responsible for hydration, $\Pi\sim 10$ kPa, is much smaller than the in-plane Young’s modulus, which is dominated by the fibre modulus $E_{f}\sim 400$ MPa. Conversely, the out-of-plane Young’s modulus is controlled by the gelatin matrix, $E_{m}\sim 4$ kPa, and is comparable to the osmotic stress; hence, there is a substantial out-of-plane deformation. When the fibre network is only partially engaged ($\lambda_{m}=3$), we see there is a larger in-plane deformation, which is due to initially inactive fibres causing a reduction in the in-plane stiffness of the material, and a smaller out-of-plane deformation. However, the in-plane stretching caused by hydration is always less than the maximum recruitment stretch. Therefore, hydration is not sufficient to fully engage the fibre network. Remarkably, the out-of-plane deformation decreases in such a way that the porosity of the material remains approximately the same in both cases. Thus, the degree of hydration is not significantly impacted by the initial level of fibre engagement.
Figure 4: Hydration of fibre-reinforced hydrogels without ($\lambda_{m}=1$, solid lines) and with ($\lambda_{m}=3$, dashed lines) fibre recruitment. The panels show the in-plane stretch (a, e), out-of-plane stretch (b, f), porosity of the hydrated state (c, g), and the osmotic stress (d, h) as a function of the Flory interaction parameter, $\chi$, with $\Phi_{f}=0.26$ (a-d), and the nominal fibre fraction, $\Phi_{f}$, with $\chi=0.57$ (e-h).
@@
By fixing $\chi=0.57$ and considering fibre networks that are initially fully engaged ($\lambda_{m}=1$), we see that the deformation and porosity do not strongly depend on the nominal fibre fraction (Fig. 4 (e)–(g)). The osmotic stress decreases linearly with $\Phi_{f}$ (Fig. 4 (h)), reflecting the balance between osmotic stress and the out-of-plane elastic stress produced by the gelatin matrix, the latter of which is proportional to $1-\Phi_{f}$; see Appendix C. When initially inactive fibres are accounted for ($\lambda_{m}=3$), hydration leads to greater in-plane stretching and reduced out-of-plane stretching compared to the active case (as is the case when these parameters are plotted as a function of $\chi$). Moreover, both the in-plane and out-of-plane stretches now strongly depend on the fibre fraction $\Phi_{f}$. The dependence of the stretches on $\Phi_{f}$ ultimately results from the in-plane softening that occurs due to inactive fibres. This softening allows the osmotic stress to drive a greater in-plane expansion of the material. As the nominal fibre fraction increases, the in-plane stiffness increases as well, reducing the in-plane deformation and increasing the out-of-plane deformation. Again, whether the fibres are initially active or inactive does not affect the porosity as a function of $\Phi_{f}$ (Fig. 4 (g)).